OpenStudy (anonymous):

THE HYPOTENUSE OF AN ISOSCELES RIGHT TRIANGLEMIS 30cm, FIND THE LENGTH OF THE OTHER SIDE >?(with medal)

OpenStudy (anonymous):

@robtobey not responding

OpenStudy (yttrium):

you know pythagoren theorem right? a^2 + b^2 = c^2 so knowing that it is isosceles, we can just let a=b hence a^2 + a^2 = c^2 2a^2 = c^2 a = sqrt(c^2/2)